OpenStudy (anonymous):

Karlo's father is 26 years older than Karlo. The sum of their ages five years ago is 42. How old are they now?

OpenStudy (anonymous):

39 and 13

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

father-39 years karlo- 13 years

OpenStudy (anonymous):

howd you come up wuth that answer?

OpenStudy (anonymous):

let x be the age of karlo so fathers age is x+26 given sum = 42 so x+(x+26)=42 2x+26=42 x=8 which is karlos age father = x+26 = 34 after 5 years just add 5 to their ages so 39 and 13

OpenStudy (anonymous):

sorry 5 years back their ages would be 8 and 34 presently their age is 13 and 39

OpenStudy (anonymous):

Let age of Karlo's father=x age of Karlo=y then x-y=26 Five years before age of Karlo's father =x-5 age of Karlo=y-5 By the question x-5+y-5=42 or x+y=42+10=52 x+y=52 x-y=26 adding 2x=78 x=39 y=52-39=13 age of Karlo=13 age of Karlo's father= 39